OpenStudy (anonymous):

what is the value of a? 9a-3(a-6)=-6

OpenStudy (uri):

9a-3a+12=-6 6a=-6-12 6a=-18 a=-18/6 a=-3

OpenStudy (anonymous):

@uri thanks